How many electrons move through a single 35.0 W headlight? How many electrons pass through it each second

1
Expert's answer
2022-03-06T15:17:37-0500

The power of a current

"P=VI"

The current

"I=\frac{P}{V}"

Hence, the number of electrons passing through the bulb

"N=\frac{q}{e}=\frac{It}{e}=\frac{Pt}{Ve}"

"N=\frac{35.0*1}{220*1.6*10^{-19}}=9.94*10^{17}"